# SLM-MUX: Orchestrating Small Language Models for Reasoning

### v1

@@ Results
The effect of “Adding More Participating Model Types” varies substantially across benchmarks. On GPQA, accuracy peaks when combining two models and declines thereafter. On GSM8K, accuracy quickly saturates at two models without further gains. In contrast, on MATH, accuracy continues to improve as additional models are included. Despite these differences, the union accuracy of model orchestration consistently increases with more models, emphasizing the role of output contradictions among models, as elaborated in Section 3.2.
@@
Figure 8: Drawing More Samples per Model Improves Accuracy. We report mean accuracy of SLM-MUX as the number of samples per model increases from 2 to 9 across three benchmarks. Group 1 and Group 2 are from Table 2. We also plotted the mean accuracy of Agent Forest (Li et al., 2024) in grey line.
Figure 9: Adding More Participating Models Affects Accuracy Differently. We report the mean accuracy (blue line) of the optimal SLM-MUX obtained when using 2 to 5 models across three benchmarks. We also report the union accuracy (grey line), defined in Section 3.2. The blue line (Mean Accuracy) is plotted against the left-hand Y-axis. The grey line (Union Accuracy) is plotted against the right-hand Y-axis.
@@ 5 Discussion
Mathematical Analysis
To explain our good performance, we note a limitation of self-consistency methods. Suppose a model has probability $p$ of answering a question correctly. When self-consistency samples $N$ responses, the probability of obtaining the correct answer after aggregation follows a binomial distribution.
$A(N,p)=\Pr\left(X\geq\left\lceil\tfrac{N}{2}\right\rceil\right)=\sum_{k=\lceil N/2\rceil}^{N}\binom{N}{k}p^{k}(1-p)^{N-k},\quad X\sim\text{Binomial}(N,p)$
$A(N,p)$ exceeds $p$ only when $p>0.5$, meaning self-consistency is effective only in this regime. When $p<0.5$, however, self-consistency can actually lower overall accuracy. For any dataset, we can conceptually divide examples into three types of questions. Type 1 includes cases where $p=100\%$, so the LLM always answers correctly. Type 2 covers cases where $p>50\%$, meaning the model is more likely than not to be correct. Type 3 includes cases where $p<50\%$, where the model is more likely to be wrong. The overall effect of self-consistency is then the improvement from Type 2, offset by the degradation from Type 3. Improvement occurs only when the dataset contains a sufficiently large proportion of Type 2 questions.
For the SLM-MUX, we select the output from the most confident model, so the accuracy can be approximated as $A(N,p_{\max})$, where $p_{\max}$ is the highest probability among the three models. By increasing $p_{\max}$, we effectively enlarge the proportion of Type 2 questions, leading to higher overall accuracy. For the Agent Forest approach, answers are drawn evenly from all models, so its accuracy can be approximated as $A(N,\bar{p})$, where $\bar{p}$ is the average probability across models. This generally results in lower accuracy than SLM-MUX.
@@ Conclusion
This work demonstrates that orchestration methods designed for frontier models paradoxically degrade the performance of SLMs by amplifying errors. To address this, we propose SLM-MUX, a framework that avoids inter-model discussion, instead selecting the most reliable output based on each model’s self-consistency. We further introduce a model selection search algorithm to find complementary model combinations. Experiments show our method not only substantially outperforms existing strategies but also enables an ensemble of just two SLMs to surpass the much larger Qwen-2.5 72B model on key reasoning benchmarks. In summary, our work validates that intelligently orchestrating multiple efficient models—a "multi-core" approach—is a highly promising alternative to endlessly scaling monolithic models on the path toward more capable AI systems.
@@ Acknowledgments
We thank Prof. Tom Griffiths (Princeton University) for insightful technical discussions, which inspired our early exploration of our method.
@@
The appendix is organized as follows. Section B presents additional visual examples illustrating the workflow and effectiveness of the SLM-MUX method across the MATH, GPQA, and GSM8K datasets. Section C provides a detailed analysis of SLM failures in discussion-based orchestration methods, drawing on experiment logs to highlight common failure patterns. Section D reports the accuracy of individual models used in our experiments. Finally, Section E provides the licensing details for the datasets.
@@ Appendix A LLM Usage Statement
Appendix B Additional Visual Illustrations of SLM-MUX
@@
Appendix C Detailed Analysis of SLM Failures in Discussion-Based Methods
@@
Appendix D Accuracy of Single LLMs
Model MATH Acc (%) GPQA Acc (%) GSM Acc (%)
Llama-3.1-8B 48.6 23.7 84.2
[…]
Gemma-2-27B 56.8 38.8 81.6
Model MATH GPQA
Accuracy (%) Token Usage Accuracy (%) Token Usage
DeepSeek V3 87.0 419,513 55.1 173,885
Gemini 2.0 Flash 90.4 361,737 63.6 195,576
GPT-4o 79.8 408,410 51.0 212,037
Appendix E Licenses for Datasets

### v2

@@ Results
The effect of “Adding More Participating Model Types” varies substantially across benchmarks. On GPQA, accuracy peaks when combining two models and declines thereafter. On GSM8K, accuracy quickly saturates at two models without further gains. In contrast, on MATH, accuracy continues to improve as additional models are included. Despite these differences, the union accuracy of model orchestration consistently increases with more models, emphasizing the role of output contradictions among models, as elaborated in Section 3.2. We also validate this scaling behavior on the test set; see Appendix D.4 for details.
@@
Figure 8: Drawing More Samples per Model Improves Accuracy. We report mean accuracy of SLM-MUX as the number of samples per model increases from 2 to 9 across three benchmarks. Group 1 and Group 2 are from Table 2. We also plot the mean accuracy of Agent Forest (Li et al., 2024) in grey line.
Figure 9: Adding More Participating Models Affects Accuracy Differently. We report the mean accuracy (blue line) of the optimal SLM-MUX obtained when using 2 to 5 models across three benchmarks. We also report the union accuracy (grey line), defined in Section 3.2. The blue line (Mean Accuracy) is plotted against the left-hand Y-axis. The grey line (Union Accuracy) is plotted against the right-hand Y-axis. For each $K$, both curves correspond to the single model combination that maximizes our model selection objective (Section 3.2) under that fixed $K$.
@@ 5 Discussion
Mathematical Intuition behind SLM-MUX
Different SLMs have complementary strengths: for any given question, some models are more likely to answer correctly than others. SLM-MUX exploits this by selecting the most self-consistent model’s output through a simple rule-based mechanism that requires no inter-model communication.
The key insight is that the confidence score can identify the strongest model. We assume that for each question, there is a unique correct answer, while incorrect answers are scattered rather than clustered. Under this assumption, a model with higher accuracy $p_{i}$ produces the correct answer more frequently across $N$ samples, leading to a higher confidence score. Therefore, selecting the model with the highest confidence score effectively identifies the model most likely to be correct.
More formally, consider $K$ models where model $i$ has probability $p_{i}$ of being correct. Let $i^{*}=\arg\max_{i}p_{i}$ denote the strongest model with margin $\gamma=p_{i^{*}}-\max_{j\neq i^{*}}p_{j}>0$. Under our assumption, the confidence score $s_{i}$ (the frequency of the most common answer over $N$ samples) concentrates around $p_{i}$. Applying Hoeffding’s inequality and a union bound, the probability of correctly selecting the strongest model satisfies:
$\Pr(\hat{i}=i^{*})\geq 1-2(K-1)\exp\left(-\frac{N\gamma^{2}}{2}\right).$
This bound shows that the probability of misidentifying the strongest model decays exponentially with sample size $N$.
This selection mechanism explains why SLM-MUX outperforms alternatives. Unlike a single fixed model, SLM-MUX performs per-question routing, effectively achieving accuracy $p_{\max}$ by always selecting the strongest available expert. Unlike pooling methods such as Agent Forest that aggregate outputs from all models, SLM-MUX avoids interference from weaker models. For instance, if the strongest model has $p_{1}=0.8$ and a weaker one has $p_{2}=0.3$, pooling their outputs merely dilutes the correct answer’s frequency. By isolating the strongest model and selecting its most frequent answer, SLM-MUX preserves the full predictive power of the most reliable source. We provide a more detailed comparative analysis with self-consistency and Agent Forest in Appendix D.2.
Extending SLM-MUX to Open-Ended Generation.
Although the current implementation of SLM-MUX relies on majority voting and is therefore restricted to tasks with discrete answer spaces, the underlying idea of selecting the most self-consistent model is more general. For open-ended generation, one can replace majority voting with alternative consistency estimators, such as LLM-as-a-judge scoring or embedding-based similarity measures. In Appendix E.1, we show a simple extension of SLM-MUX to HumanEval (Chen et al., 2021) using this idea and observe strong empirical gains.
Extending SLM-MUX Beyond Generalist SLMs.
The experiments above focus on general-purpose SLMs. We further evaluate whether the consistency-based selection principle extends to other settings: (1) frontier LLMs such as GPT-4o and Gemini-2.5-Flash, and (2) domain-specific fine-tuned models such as code and math specialists. In both cases, SLM-MUX achieves consistent improvements over the best single model. Full experimental details are provided in Appendix E.
@@ Conclusion
This work demonstrates that orchestration methods designed for frontier models paradoxically degrade the performance of SLMs by amplifying errors. To address this, we propose SLM-MUX, a framework that avoids inter-model discussion, instead selecting the most reliable output based on each model’s self-consistency. We further introduce a model selection search algorithm to find complementary model combinations. Experiments show our method not only substantially outperforms existing strategies but also enables an ensemble of just two SLMs to surpass the much larger Qwen-2.5 72B model on key reasoning benchmarks. In summary, our work validates that intelligently orchestrating multiple efficient models—a "multi-core" approach—is a promising alternative to scaling monolithic models on the path toward more capable AI systems.
@@ Acknowledgments
We thank Tom Griffiths from Princeton University, whose insights and feedback were instrumental in initiating this project.
@@
The appendix is organized into five main sections. Section A states the usage of LLMs in preparing this paper. Section B provides experimental details, including visual illustrations, single-model accuracies, and standard deviation calculations. Section C analyzes why discussion-based methods fail on SLMs, presenting groupthink analysis and prompt sensitivity studies. Section D validates the SLM-MUX design through consistency-accuracy correlation analysis, comparative analysis with voting-based methods, model selection search analysis, and test-set scaling validation. Section E demonstrates the generalization of SLM-MUX to open-ended generation, frontier LLMs, and domain-specific models. Finally, Section F provides dataset licenses.
@@ Appendix A LLM Usage Statement
Appendix B Experimental Details
B.1 Visual Illustrations of SLM-MUX
@@
B.2 Accuracy of Single LLMs
Model MATH Acc (%) GPQA Acc (%) GSM Acc (%)
Llama-3.1-8B 48.6 23.7 84.2
[…]
Gemma-2-27B 56.8 38.8 81.6
Model MATH GPQA
Accuracy (%) Token Usage Accuracy (%) Token Usage
DeepSeek V3 87.0 419,513 55.1 173,885
Gemini 2.0 Flash 90.4 361,737 63.6 195,576
GPT-4o 79.8 408,410 51.0 212,037
B.3 Standard Deviation of the Data Points in Figure 5
Although all experiments are run in a deterministic setting with temperature set to zero, we can still compute the standard deviation of each datapoint by treating the outcome as a Bernoulli variable. Specifically, if there are $n_{\text{correct}}$ correct answers and $n_{\text{wrong}}$ incorrect answers, the standard deviation is
$\frac{\sqrt{\operatorname{Var}(X)}}{\sqrt{n_{\text{total}}}}=\frac{\sqrt{p(1-p)}}{\sqrt{n_{\text{total}}}}=\frac{\displaystyle\sqrt{\frac{n_{\text{correct}}}{n_{\text{total}}}\left(1-\frac{n_{\text{correct}}}{n_{\text{total}}}\right)}}{\sqrt{n_{\text{total}}}},$
where $n_{\text{total}}=n_{\text{correct}}+n_{\text{wrong}}$.
The results are summarized in Table 7.
Table 7: Accuracy and estimated standard deviation on MATH ($n=500$) and GPQA ($n=196$) using datapoints from Figure 5.
MATH ($n=500$) GPQA ($n=196$)
Method SLM orchestration LLM composition SLM orchestration LLM composition
Single Model Max $56.8\pm 2.22$ $90.4\pm 1.32$ $38.8\pm 3.48$ $63.6\pm 3.44$
Mixture-of-Agents $46.2\pm 2.23$ $88.8\pm 1.41$ $33.3\pm 3.37$ $58.6\pm 3.52$
LLM-Debate $51.6\pm 2.23$ $90.8\pm 1.29$ $36.9\pm 3.45$ $65.6\pm 3.39$
Multi-Agent Verification $48.4\pm 2.23$ $91.6\pm 1.24$ $35.4\pm 3.42$ $64.2\pm 3.42$
Appendix C Why Discussion-Based Methods Fail on SLMs
C.1 Groupthink Analysis
@@
C.2 Prompt Sensitivity Analysis
A natural concern is whether the performance gap between discussion-based methods and SLM-MUX is due to suboptimal prompt selection rather than inherent limitations. To address this, we conduct a prompt sensitivity analysis on the Mixture-of-Agents (MoA) baseline using the MATH benchmark.
We use Gemini-2.5-Flash to generate 10 diverse aggregator prompts for MoA. Table 9 summarizes the results. Our baseline prompt (46.2%) outperforms the average of the tuned prompts (41.8%) and surpasses 7 out of 10 generated prompts. Even the best tuned prompt (48.4%) remains substantially below the best single model (56.8%) and SLM-MUX (61.8%).
Table 9: Prompt sensitivity analysis for MoA on MATH. The baseline prompt used in our experiments is already near-optimal.
Setting Accuracy
SLM-MUX (Ours) 61.8%
Best Single Model 56.8%
Tuned Prompt (Best) 48.4%
Baseline Prompt (Used in Paper) 46.2%
Tuned Prompt (Average) 41.8%
Tuned Prompt (Worst) 25.8%
To further stress-test this result, we conducted iterative prompt optimization directly on the test set for 6 rounds, selecting the best-performing prompt at each iteration. Even under this extremely favorable setting for MoA, the accuracy peaked at 55.6%, still below the best single model (56.8%) and far below SLM-MUX (61.8%). This confirms that the performance gap reflects inherent limitations of discussion-based aggregation when applied to SLMs, rather than an artifact of prompt selection.
Appendix D Validation of SLM-MUX Design
D.1 Consistency vs Accuracy Correlation
We empirically study how per-question self-consistency correlates with accuracy on four datasets: GSM8K, MATH, GPQA, and HumanEval. For each model–dataset pair, we compute a self-consistency score for every question (as defined in the main text) and group questions into three bins according to this score: Low $[0.0,0.5)$, Medium $[0.5,0.8)$, and High $[0.8,1.0]$. We then measure the empirical accuracy (fraction of correct answers) within each bin.
Figure 15 reports the resulting accuracies for two representative SLMs on each dataset. Across GSM8K, MATH, and HumanEval we observe a strong positive relationship between self-consistency and accuracy: questions in the high-consistency bin are substantially more likely to be answered correctly than those in the low-consistency bin. GPQA exhibits a weaker but still positive correlation. Overall, these results provide empirical support for the link between self-consistency and correctness assumed in our method.
Figure 15: Accuracy as a function of per-question self-consistency score on GSM8K, MATH, GPQA, and HumanEval. Questions are grouped into three bins by self-consistency: Low $[0.0,0.5)$, Medium $[0.5,0.8)$, and High $[0.8,1.0]$. Each line corresponds to a different SLM; legends report the Pearson correlation $r$ between self-consistency and correctness.
D.2 Comparative Analysis with Voting-Based Methods
Since SLM-MUX also involves voting on model outputs, we examine its differences from standard self-consistency and Agent Forest to better explain the source of our improvements.
To explain our stronger performance, we note a limitation of self-consistency methods. Suppose a model has probability $p$ of answering a question correctly. When self-consistency samples $N$ responses, the probability of obtaining the correct answer after majority voting follows a binomial distribution:
$A(N,p)=\Pr\left(X\geq\left\lceil\tfrac{N}{2}\right\rceil\right)=\sum_{k=\lceil N/2\rceil}^{N}\binom{N}{k}p^{k}(1-p)^{N-k},\quad X\sim\text{Binomial}(N,p)$ (1)
We observe that $A(N,p)$ exceeds $p$ only when $p>0.5$, meaning self-consistency is effective only in this regime. When $p<0.5$, self-consistency can actually lower overall accuracy.
For any dataset, we can conceptually divide examples into three types of questions. Type 1: $p=100\%$, the model always answers correctly. Type 2: $p>50\%$, the model is more likely than not to be correct. Type 3: $p<50\%$, the model is more likely to be wrong. The overall effect of self-consistency is then the improvement from Type 2, offset by the degradation from Type 3. Improvement occurs only when the dataset contains a sufficiently large proportion of Type 2 questions.
For SLM-MUX, we select the output from the most confident model, so the accuracy can be approximated as $A(N,p_{\max})$, where $p_{\max}$ is the highest probability among the participating models. By routing to the model with the highest $p_{\max}$ on each question, we effectively enlarge the proportion of Type 2 questions, leading to higher overall accuracy.
For the Agent Forest approach, answers are drawn evenly from all models, so its accuracy can be approximated as $A(N,\bar{p})$, where $\bar{p}$ is the average probability across models. This generally results in lower accuracy than SLM-MUX, as weaker models dilute the signal from stronger ones.
D.3 Model Selection Search Analysis
D.3.1 Search-Set Size Stability
We analyze how the size of the search set (number of problems used in the search phase) affects the resulting model ranking. For each of the three benchmarks MATH, GSM8K, and GPQA, we treat the full benchmark (approximately 500 problems) as the search pool and first run our search procedure on the full set to obtain a ranking of all candidate models. We then record the models occupying Rank 1, Rank 2, and Rank 3 under this full-set ranking.
Next, we subsample the search set to smaller sizes and re-run the search. Specifically, for each dataset we evaluate search-set sizes of $100$, $200$, $300$, $400$, and the full set. At each size, we recompute the ranking over all models and track the ranks of the three models that were Rank 1–3 under the full-set setting.
Figure 16 shows the results. Each curve corresponds to one of the Rank 1/2/3 models under the full-set ranking, and the $y$-axis reports its rank when the search-set size is changed. Across all three datasets, these models consistently remain within the top three positions even when the search set is reduced to as few as 100 problems, with only minor swaps in their relative order on MATH and GSM8K. This suggests that a search set of a few hundred problems is sufficient to stably identify the top-performing models.
Figure 16: Stability of model rankings with respect to search-set size on MATH, GSM8K, and GPQA. For each dataset, we first determine the top three models using the full search set and then track their ranks as the search-set size is reduced.
D.3.2 Random Model Selection Baseline
A natural question is whether the performance gains of SLM-MUX stem from the model selection search or from the orchestration architecture itself. To isolate the contribution of the architecture, we conduct an experiment where model combinations are selected randomly rather than through our search procedure.
For each dataset, we randomly sample model combinations of size $K=2,3,4$ from our pool of five SLMs and apply SLM-MUX. We compare the resulting accuracy against the best single model within each random pool. Table 10 reports the average performance across all random combinations.
Table 10: Performance of SLM-MUX with randomly selected model combinations. Even without optimized model selection, SLM-MUX consistently outperforms the best single model in the pool.
Dataset $K$ SLM-MUX (Random) Best Single Model $\Delta$ (Gain)
MATH-500 2 67.2% 66.1% +1.1
3 71.5% 69.8% +1.7
4 73.7% 71.8% +1.9
GSM8K 2 84.1% 81.7% +2.4
3 87.3% 83.6% +3.7
4 88.8% 84.0% +4.8
GPQA 2 43.0% 41.3% +1.7
3 44.5% 44.4% +0.1
4 46.3% 46.0% +0.3
As shown in Table 10, SLM-MUX consistently outperforms the best single model even when the model combination is selected randomly. On MATH-500 and GSM8K, the gains are substantial (up to +4.8 on GSM8K with $K=4$). On GPQA, where consistency is a weaker signal for correctness, the gains are smaller but still positive.
To further quantify how often SLM-MUX improves over single models, we compute the Effective Combination Rate: the percentage of all possible $K$-model subsets where SLM-MUX outperforms the best single model in that subset. Table 11 reports the results.
Table 11: Effective Combination Rate: percentage of model combinations where SLM-MUX outperforms the best single model in the subset.
Dataset $K=2$ $K=3$ $K=4$
MATH-500 70% 100% 100%
GSM8K 100% 100% 100%
GPQA 60% 80% 100%
On GSM8K, 100% of combinations are effective across all values of $K$. On MATH-500 and GPQA, the effective rate increases with $K$, reaching 100% at $K=4$. Even at $K=2$, the majority of combinations (60–100%) are effective. These results demonstrate that the space of “workable” model combinations is dense, and one does not need to search extensively to find an effective subset. The model selection search provides additional gains by identifying the optimal combination, but the architecture is robust even without it.
D.4 Test-Set Validation of Scaling Behavior
In Section 3.3, we evaluated the “Adding More Participating Model Types” scaling dimension on the validation set. Here we report the corresponding results on the test set to verify that the observed trends generalize.
Table 12 summarizes the test-set accuracy as the number of participating models $K$ increases from 1 to 5. For each value of $K$, we use the model combination selected by the search procedure described in Section 3.2. Figure 17 visualizes these results.
Table 12: Test-set accuracy (%) as the number of participating models $K$ increases. For $K=1$, the best single model is reported.
$K$ GPQA GSM8K MATH
1 (best single) 47.98 90.20 78.20
2 50.21 91.51 80.27
3 51.81 91.29 81.09
4 49.16 91.44 81.00
5 45.18 90.80 80.64
Figure 17: Test-set accuracy as a function of the number of participating models on GPQA, GSM8K, and MATH. The trends mirror those observed on the validation set: GPQA peaks at $K=3$, GSM8K saturates quickly, and MATH shows continued improvement up to $K=3$.
Appendix E Generalization of SLM-MUX
E.1 Open-Ended Generation (HumanEval)
In the main text, SLM-MUX is instantiated on tasks with discrete answer spaces, where self-consistency can be measured via majority voting over sampled outputs. To check whether the same principle extends to open-ended generation, we apply SLM-MUX to the HumanEval code-generation benchmark.
Consistency estimator for open-ended generation.
On HumanEval, exact-string majority voting is not appropriate, so we replace it with a semantic consistency estimator. For each model and each problem, we sample $N=5$ code generations with temperature $0.3$. We then encode the 5 generations using the pretrained embedding model Salesforce/codet5p-110m-embedding and compute pairwise cosine similarities, yielding a $5\times 5$ similarity matrix. From this matrix, we identify the most coherent cluster of generations (of size $x$) and use $x/5$ as the model’s self-consistency score for that problem, analogous to the confidence score derived from majority voting in the discrete-answer setting.
We also experimented with an LLM-as-a-judge–based consistency estimator (using Qwen2.5-7B-Instruct as the judge) and found that the embedding-based estimator exhibits a stronger correlation with ground-truth correctness (Pass@1). All results below therefore use the embedding-based consistency score.
Results.
Table 13 reports the Pass@1 of each individual SLM on HumanEval. Table 14 reports the Pass@1 of SLM-MUX when combining two models at a time; for each problem, SLM-MUX selects the solution from the model with the larger embedding-based consistency score.
Table 13: Pass@1 on HumanEval for individual SLMs.
Model Pass@1
Llama-3.1-8B-Instruct 0.178
Qwen2.5-7B-Instruct 0.485
Mistral-Small-24B 0.870
Qwen2.5-Coder-7B 0.893
Table 14: SLM-MUX on HumanEval (Pass@1). Each row corresponds to a pair of SLMs; SLM-MUX selects the output from the model with higher embedding-based consistency.
Setup Models combined SLM-MUX (Pass@1)
Exp 1 Llama-3.1-8B-Instruct + Qwen2.5-7B-Instruct 0.506
Exp 2 Mistral-Small-24B + Qwen2.5-Coder-7B 0.939
E.2 Frontier LLMs
We evaluate whether SLM-MUX can exploit complementary strengths between state-of-the-art frontier models. We pair GPT-4o with Gemini-2.5-Flash and apply SLM-MUX on MATH-500, GPQA, and GSM8K-500. For each problem, we sample $N=5$ responses per model at temperature 0.3 and apply self-consistency routing: for each problem, we perform majority voting within each model’s samples, then route to the model showing higher agreement.
Table 15 summarizes the results. As a reference, “Perfect Routing” indicates the theoretical upper bound achievable if the system always selects the correct model when at least one succeeds.
Table 15: SLM-MUX performance when applied to frontier LLMs (GPT-4o and Gemini-2.5-Flash).
Benchmark GPT-4o Gemini-2.5-Flash SLM-MUX Perfect Routing
MATH-500 73.0% 92.1% 92.8% 94.2%
GPQA 50.7% 51.1% 60.1% 73.7%
GSM8K-500 89.1% 85.7% 89.4% 91.4%
The results reveal two distinct regimes. On GPQA, the two models exhibit complementary error patterns, and SLM-MUX achieves 60.1% accuracy, surpassing the best single model by nearly 10 percentage points. This demonstrates that SLM-MUX effectively exploits complementary strengths even at the frontier scale. On MATH and GSM8K, the Perfect Routing bounds (94.2% and 91.4%) are only marginally higher than the single-model baselines, indicating high overlap in the models’ correct predictions. The limited gains on these benchmarks reflect this ceiling rather than a limitation of the routing mechanism.
E.3 Domain-Specific Fine-Tuned Models
Domain-specific fine-tuned models are widely deployed in practice. We evaluate whether SLM-MUX can effectively orchestrate such specialized models by testing on two domains: code generation and mathematical reasoning.
For code generation, we pair Qwen2.5-Coder-7B (a code-specialized model) with Mistral-Small-24B (a general-purpose model) on HumanEval. We use the same embedding-based consistency estimator described in Section E.1. For mathematical reasoning, we pair DeepSeek-Math-7B-RL (a math-specialized model) with Llama-3.1-8B-Instruct on MATH-500, using standard majority voting for consistency estimation. Both experiments sample $N=5$ responses per model at temperature 0.3.
Table 16: SLM-MUX performance when orchestrating domain-specific fine-tuned models.
Domain Benchmark Best Single Model SLM-MUX $\Delta$ (Gain)
Code Generation HumanEval 89.3% 93.9% +4.6
Math Reasoning MATH-500 58.8% 62.2% +3.4
As shown in Table 16, SLM-MUX achieves consistent improvements in both domains. On HumanEval, the orchestrated system reaches 93.9% Pass@1, outperforming the code specialist (89.3%) by 4.6 percentage points. On MATH-500, combining the math specialist with a general-purpose model yields 62.2% accuracy, a 3.4pp improvement. These results demonstrate that SLM-MUX generalizes effectively to domain-specific fine-tuned models, successfully capturing complementary strengths between specialists and generalists.
Appendix F Licenses for Datasets